# Random pairing MLE for estimation of item parameters in Rasch model

### v1

@@ E.1 Proof of Theorem 3
$-(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{B}\widehat{\bm{\epsilon}}$ $\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right).$
As $m$ is fixed, Lemma 2 implies that $\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}\rightarrow\infty$ is equivalent to $np\rightarrow\infty$. For the residual term, Lemma 4 and Lemma 3 imply that with probability at least $1-O(n^{-10})$,
$\|\bm{L}_{Lz}^{1/2}\bm{r}\|\leq\|\bm{L}_{Lz}^{1/2}\|\cdot\|\bm{r}\|\leq\|\bm{L}_{Lz}^{1/2}\|\cdot\sqrt{m}\|\bm{r}\|_{\infty}\leq\frac{C_{1}\sqrt{m}\kappa_{1}^{6}\log^{2}(n)}{\sqrt{np}}.$
Then as $np/\log^{4}(n)\rightarrow\infty$,
$\bm{L}_{Lz}^{1/2}\bm{r}\rightarrow\bm{0}$
in probability. Combining the limiting distribution of both two terms, we have
@@
$\bm{L}_{Lz}^{1/2}\widehat{\bm{\delta}}=-(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{B}\widehat{\bm{\epsilon}}+\bm{L}_{Lz}^{1/2}\bm{r}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right).$
In the case of $\widetilde{\bm{z}}=\widehat{\bm{z}}$, we introduce the following lemma. The proof is deferred to the end of this section.
@@
Instate the assumptions of Theorem 1. Then with probability at least $1-C_{1}n^{-10}$ for some constant $C_{1}$, one has
$\|\bm{L}_{Lz}^{1/2}-\bm{L}_{L\hat{z}}^{1/2}\|\leq C_{2}\kappa_{1}^{2}\kappa_{2}^{1/2}\sqrt{\log(n)}$
for some constant $C_{2}$.
Now the above lemma and Theorem 1 imply
$\left\|\bm{L}_{L\widehat{z}}^{1/2}\widehat{\bm{\delta}}-\bm{L}_{Lz}^{1/2}\widehat{\bm{\delta}}\right\|\leq\|\bm{L}_{L\widetilde{z}}^{1/2}-\bm{L}_{Lz}^{1/2}\|\cdot\|\widehat{\bm{\delta}}\|\leq C_{3}\kappa_{1}^{3}\kappa_{2}\sqrt{\frac{m\log^{2}(n)}{np}}$
for some constant $C_{3}>0$. Therefore, as $np/\log^{4}(n)\rightarrow\infty$,
$\bm{L}_{L\widehat{z}}^{1/2}\widehat{\bm{\delta}}-\bm{L}_{Lz}^{1/2}\widehat{\bm{\delta}}\rightarrow 0$
in probability and hence
$\bm{L}_{L\widehat{z}}^{1/2}\widehat{\bm{\delta}}=\left(\bm{L}_{L\widehat{z}}^{1/2}\widehat{\bm{\delta}}-\bm{L}_{Lz}^{1/2}\widehat{\bm{\delta}}\right)+\bm{L}_{Lz}^{1/2}\widehat{\bm{\delta}}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right).$
@@
Let $\bm{R}\in\mathbb{R}^{(n-1)\times n}$ be a partial isometry such that $\bm{R}\bm{R}^{\top}=\bm{I}_{m-1}$ and $\bm{R}\bm{1}=\bm{0}$. Then
$\|\bm{L}_{L\hat{z}}^{1/2}-\bm{L}_{Lz}^{1/2}\|=\|\bm{R}(\bm{L}_{L\hat{z}}^{1/2}-\bm{L}_{Lz}^{1/2})\bm{R}^{\top}\|.$
We also observe that $(\bm{R}\bm{L}_{L\widetilde{z}}\bm{R}^{\top})^{1/2}=\bm{R}\bm{L}_{L\widetilde{z}}^{1/2}\bm{R}^{\top}$ for $\widetilde{z}\in\{z,\widehat{z}\}$. As shown in the proof of Lemma 7,
$\|\bm{L}_{L\hat{z}}-\bm{L}_{Lz}\|\leq C_{1}\kappa_{1}^{3/2}\sqrt{np\log(n)}$
for some constant $C_{1}>0$.
Lemma 3 implies that
$\lambda_{m-1}(\bm{R}\bm{L}_{Lz}^{1/2}\bm{R}^{\top})=\left[\lambda_{m-1}(\bm{L}_{Lz})\right]{}^{1/2}\geq\frac{\sqrt{np}}{4\kappa_{1}^{1/2}\kappa_{2}^{1/2}}.$
Then by Lemma 2.1 in [Sch92],
$\|\bm{R}(\bm{L}_{L\widehat{z}}^{1/2}-\bm{L}_{Lz}^{1/2})\bm{R}^{\top}\|\leq\frac{\|\bm{R}(\bm{L}_{L\widehat{z}}-\bm{L}_{Lz})\bm{R}^{\top}\|}{\lambda_{m-1}(\bm{R}\bm{L}_{Lz}^{1/2}\bm{R}^{\top})}\leq 4C_{1}\kappa_{1}^{2}\kappa_{2}^{1/2}\sqrt{\log(n)}.$
E.2 Proof of Corollary 1
Observe that
$\frac{\widehat{\theta}_{i}-\theta_{i}^{\star}}{\sqrt{[\bm{L}_{L\tilde{z}}^{\dagger}]_{ii}}}=\frac{\bm{e}_{i}^{\top}(\bm{L}_{L\tilde{z}}^{\dagger})^{1/2}\bm{L}_{L\tilde{z}}^{1/2}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})}{\sqrt{[\bm{L}_{L\tilde{z}}^{\dagger}]_{ii}}}.$
Then Theorem 3 implies
$\frac{\bm{e}_{i}^{\top}(\bm{L}_{L\tilde{z}}^{\dagger})^{1/2}(\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top})(\bm{L}_{L\tilde{z}}^{\dagger})^{1/2}\bm{e}_{i}}{[\bm{L}_{L\tilde{z}}^{\dagger}]_{ii}}=1.$
The last inequality follows from the fact that $\bm{1}_{m}^{\top}(\bm{L}_{L\tilde{z}}^{\dagger})^{1/2}=0$. As a side note, we compare this with a similar result for the BTL model. Proposition 4.1 in [GSZ23] studies the uncertainty quantification of MLE in the BTL model with uniform observation. It says that

### v2

@@ D.1 Proof of Theorem 3
$\frac{1}{n}\sum_{t=1}^{n}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})$ $=\frac{1}{n}\sum_{t=1}^{n}\nabla\mathcal{L}_{k}^{(t)}(\widehat{\bm{\theta}}_{k}^{(n)})+\left[\int_{\tau=0}^{1}\frac{1}{n}\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star}+\tau(\widehat{\bm{\theta}}_{k}^{(n)}-\bm{\theta}^{\star}))\mathrm{d}\tau\right](\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{k}^{(n)})$
@@
$=\left[\int_{\tau=0}^{1}\frac{1}{n}\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star}+\tau(\widehat{\bm{\theta}}_{k}^{(n)}-\bm{\theta}^{\star}))\mathrm{d}\tau\right](\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{k}^{(n)}).$ (41)
Here the second line comes from the optimality condition for $\widehat{\bm{\theta}}_{k}^{(n)}$.
We first consider the Hessian part. We start by showing the almost surely convergence of $\widehat{\bm{\theta}}_{k}^{(n)}$ and controlling the $\ell_{\infty}$ norm of $\widehat{\bm{\theta}}_{k}^{(n)}$. By Theorem 1, we know that with probability at least $1-O(n^{-10})$, $\|\widehat{\bm{\theta}}_{k}^{(n)}-\bm{\theta}^{\star}\|\leq C\kappa_{1}\kappa_{2}^{1/2}\sqrt{\frac{m\log(n)}{np}}$ for some constant $C$. This and the Borel-Cantelli lemma implies that $\widehat{\bm{\theta}}_{k}^{(n)}\overset{\mathrm{a.s.}}{\rightarrow}\bm{\theta}^{\star}$ and $\|\widehat{\bm{\theta}}_{k}^{(n)}\|_{\infty}\leq 2\kappa$ for large enough $n$. Furthermore, observe that the Hessian
$\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta})=\sum_{\begin{subarray}{c}(i,j):i>j\\ (i,j,t)\in\Omega_{k}\end{subarray}}\frac{e^{\theta_{i}}e^{\theta_{j}}}{(e^{\theta_{i}}+e^{\theta_{j}})^{2}}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$
is Lipschitz-continuous for $\{\bm{\theta}:\|\bm{\theta}\|_{\infty}\leq 2\kappa\}$ with a uniform Lipschitz constant for all $k,t$. Then we conclude that
$\int_{\tau=0}^{1}\frac{1}{n}\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star}+\tau(\widehat{\bm{\theta}}_{k}^{(n)}-\bm{\theta}^{\star}))\mathrm{d}\tau\overset{\mathrm{a.s.}}{\rightarrow}\frac{1}{n}\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star}).$ (42)
Recall that the sampling, random pairing, and comparisons are all independent between different users. Moreover, the Hessian and its second moment are both bounded. Then we can invoke the strong law of large number to say that
$\frac{1}{n}\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})\overset{\mathrm{a.s.}}{\rightarrow}\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})=\bm{H}^{\infty},$ (43)
where $\bm{H}^{\infty}$ is defined in (8a). Note that this equality holds trivially for $k=1$ and then by symmetry holds all $k$. We also make the following claim. The proof is deferred to the end of this section.
@@
Instate the assumptions of Theorem 1, we have that $\lambda_{m-1}\left(\bm{H}_{k}^{\infty}\right)>0$ and $\bm{H}_{k}^{\infty\top}\bm{1}_{m}=\bm{0}_{m}$.
Now we move to the gradient part. Recall that the gradient
$\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta})$
$=\mathbb{E}_{\mathrm{s+r+d+c}}\sum_{\begin{subarray}{c}(i,j):i>j\\ (i,j,t)\in\Omega_{k}\end{subarray}}\left[\left(-\mathds{1}\{X_{ti}<X_{tj}\}+\frac{e^{\theta_{j}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)(\bm{e}_{i}-\bm{e}_{j})\mathds{1}\{X_{ti}\neq X_{tj}\}\right]$
$=\bm{0}_{m}$
is a zero-mean random variable and independent across different $t$. Recall that for all $k_{1}\in[n_{\mathrm{split}}]$,
$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{k_{1}}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{k_{1}}^{(t)}(\bm{\theta}^{\star})^{\top}=\bm{V}_{\mathrm{same}}^{\infty}.$
Note that this equality holds trivially for $k=1$ and then by symmetry holds all $k$. Similarly for all $k_{1}\neq k_{2}$,
$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{k_{1}}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{k_{2}}^{(t)}(\bm{\theta}^{\star})^{\top}=\bm{V}_{\mathrm{diff}}^{\infty}.$
Note that $\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})$ is bounded for all $k,t$. We can then invoke the central limit theorem for triangular arrays (see Theorems 3.4.10 and 3.10.6 in [12]) to reach that
$\frac{1}{\sqrt{n}}\sum_{t=1}^{n}\left[\frac{1}{n_{\mathrm{split}}}\sum_{k=1}^{n_{\mathrm{split}}}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})\right]\overset{\mathrm{d}}{\rightarrow}$
$\hskip 100.00015pt\mathcal{N}\left(\bm{0},\frac{1}{n_{\mathrm{split}}^{2}}\sum_{k_{1}=1}^{n_{\mathrm{split}}}\sum_{k_{2}=1}^{n_{\mathrm{split}}}\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{k_{1}}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{k_{2}}^{(t)}(\bm{\theta}^{\star})^{\top}\right).$
We can then rewrite the above formula as
$\frac{1}{\sqrt{n}}\sum_{t=1}^{n}\left[\frac{1}{n_{\mathrm{split}}}\sum_{k=1}^{n_{\mathrm{split}}}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})\right]\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(\bm{0},\frac{1}{n_{\mathrm{split}}}\bm{V}_{\mathrm{same}}^{\infty}+\frac{n_{\mathrm{split}}-1}{n_{\mathrm{split}}}\bm{V}_{\mathrm{diff}}^{\infty}\right).$ (44)
We now combine the gradient and the Hessian parts together. By (41), (42) and (43), we know that
$\frac{1}{n}\sum_{t=1}^{n}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})-\bm{H}^{\infty}(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{k}^{(n)})\overset{\mathrm{a.s.}}{\rightarrow}\bm{0}.$
Recall that $\bm{\theta}^{\star\top}\bm{1}_{m}=\widehat{\bm{\theta}}_{k}^{(n)\top}\bm{1}_{m}=0$ by design and $\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})^{\top}\bm{1}_{m}=0$ by definition. Combining this with Lemma 7, we can take the pseudo-inverse of $\bm{H}^{\infty}$ to reach
$\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{k}^{(n)}\overset{\mathrm{a.s.}}{\rightarrow}\left(\bm{H}^{\infty}\right)^{\dagger}\left[\frac{1}{n}\sum_{t=1}^{n}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})\right].$
Now we average across all random splittings to conclude that
$\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}-\bm{\theta}^{\star}\overset{\mathrm{a.s.}}{\rightarrow}-\left(\bm{H}^{\infty}\right)^{\dagger}\left[\frac{1}{n}\sum_{t=1}^{n}\frac{1}{n_{\mathrm{split}}}\sum_{k=1}^{n_{\mathrm{split}}}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})\right].$
Combining this with (44), we have that
$\sqrt{n}\left(\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}-\bm{\theta}^{\star}\right)\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(\bm{0},\left(\bm{H}^{\infty}\right)^{\dagger}\left[\frac{1}{n_{\mathrm{split}}}\bm{V}_{\mathrm{same}}^{\infty}+\frac{n_{\mathrm{split}}-1}{n_{\mathrm{split}}}\bm{V}_{\mathrm{diff}}^{\infty}\right]\left(\bm{H}^{\infty}\right)^{\dagger}\right).$
@@
It suffices to show that $\lambda_{m-1}(\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star}))\geq\beta n$ for some $\beta>0$ that does not depend on $n$. Recall that
$\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})$ $=\sum_{t=1}^{n}\sum_{\begin{subarray}{c}(i,j):i>j\\ (i,j,t)\in\Omega_{k}\end{subarray}}\left[\left(-\frac{e^{\theta_{i}^{\star}+\theta_{j}^{\star}}}{(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})^{2}}\right)(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\mathds{1}\{X_{ti}\neq X_{tj}\}\right]$
$=\sum_{t=1}^{n}\sum_{\begin{subarray}{c}(i,j):i>j\\ (i,j,t)\in\Omega_{k}\end{subarray}}\left[-z_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\mathds{1}\{X_{ti}\neq X_{tj}\}\right].$
Invoking Lemma 3, we have that with probability at least $1-O(n^{-10})$, $\lambda_{m-1}(\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}))\geq np/(16\kappa_{1}\kappa_{2})$, so we are done.
D.2 Proof for Theorem 4
The proof for the asymptotic normality of $\mathrm{WP\text{-}MLE}$ is very similar to the proof for $\mathrm{MRP\text{-}MLE}$. We start with a lemma on the asymptotic consistency of $\mathrm{WP\text{-}MLE}$.
Lemma 8.
Instate the assumptions of Theorem 1. Then as $n\rightarrow\infty$,
$\widehat{\bm{\theta}}_{\mathrm{WP}}^{(n)}\overset{\mathrm{p}}{\rightarrow}\bm{\theta}^{\star}.$
Recall that in the assumption of Theorem 3, we do not specified the limit of Hessian and gradient of the loss function of the $\mathrm{WP\text{-}MLE}$. In the following lemma, we show that they are implied by the limit of Hessian and gradient of the $\mathrm{MRP\text{-}MLE}$ loss function.
Lemma 9.
Let $\bm{H}^{\infty}$ and $\bm{V}_{\mathrm{diff}}^{\infty}$ be defined as in Theorem 3. Then we have that